# Dual Language Models: Balancing Training Efficiency and Overfitting Resilience

### v1

@@ 5.1 Pretraining setup
We train 470-million-parameter language models (with 360M non-embedding weights) on 32 billion tokens. This token budget is more than $4\times$ past the Chinchilla compute-optimal point (10.5555/3600270.3602446); we specifically decided to conduct the experiments in this regime as it reflects how modern language models are trained in practice. This compute budget is also large enough to induce non-trivial zero-shot downstream performance, enabling us to measure clear differences between different configurations.
@@
The language models have 24 layers with hidden size of 1 024, their self-attention operations are divided into 16 parallel heads, the feed-forward modules have intermediate size of 3 554, and the vocabulary is set to 51 200 tokens. As for the architecture itself, we follow the usual modifications of the original transformer recipe (NIPS2017_3f5ee243) – pre-normalization (nguyen-salazar-2019-transformers) with RMSNorm (10.5555/3454287.3455397), rotational positional embedding (10.1016/j.neucom.2023.127063) and Swish-gated linear units (ramachandran2018searching; shazeer2020gluvariantsimprovetransformer).
@@
The parameters are optimized by the Muon optimizer for faster convergence (jordan2024muon), specifically its variation proposed by liu2025muonscalablellmtraining. The learning rate is set to 0.007 and decayed according to the warmup-stable-decay (WSD; hagele2024scaling) schedule (without warmup steps and 2 048 steps of linear decay). In total, each model is trained for 8 192 steps with 4M tokens in each global batch and with a sequence length of 2 048 tokens. The optimization is regularized by weight decay (with strength of $10^{-1}$) and by an auxiliary z-loss term (with strength of $10^{-4}$; chowdhery2022palmscalinglanguagemodeling).
@@
Even though we limit the training data to 32B tokens, we deliberately choose a text corpus that is not excessively filtered and that is representative of large-scale web crawls used in practice. We randomly sample English documents with 32B tokens in total from the HPLT v2 corpus (burchell-etal-2025-expanded), which combines extracted webpages from the Internet Archive and CommonCrawl. We also use a smaller disjoint subset to monitor the validation loss. To prevent a potential bias from using an external tokenizer, we train a standard byte-level BPE tokenizer (10.5555/177910.177914) with 51 200 subwords directly on the full training data.
5.2 Finding the optimal autoregressive-diffusion ratio
We trained and evaluated 50 language models in total, the results are plotted in Figure 4. In order to deal with the noisy nature of this data and to better understand the relation between the amount of data repetitions and the optimal autoregressive-diffusion ratio, we use simple statistical models.
@@ 5.2 Finding the optimal autoregressive-diffusion ratio
We use Gaussian process regression (GPR; NIPS1995_7cce53cf) with a composite kernel structure to model the relationship between data repetitions, AR-D ratios and downstream performance. The kernel consists of a constant kernel multiplied by an anisotropic Matérn kernel ($\nu=1.5$; stein1999interpolation) combined additively with a white noise kernel to account for observation noise. The input features are standardized to zero mean and unit variance, and the output features are normalized. The kernel parameters are optimized by L-BFGS-B (10.5555/3112655.3112866) using SciPy (2020SciPy-NMeth). The resulting interpolations in Figure 4 show regular structure while closely fitting the data with $R^{2}$ over 0.99 in all cases.
Figure 4: Interpolated unidirectional and bidirectional results. The (a) and (b) figures on top show the relation between repetitions (x-axis) and autoregressive-diffusion ratios (y-axis); the contours follow the Gaussian process model that interpolates the average performance of language models trained according to the specified settings. The respective results are plotted either as crosses when the model overfitted during training, or as circles. The (c) and (d) figures below visualize the estimated probability that a particular ratio (y-axis) is optimal for a given number of repetitions (x-axis).
@@
The fitted Gaussian process is a probabilistic model of the downstream performance with regard to the amount of data repetition and the AR-D ratio. Thus, we can transform this to the probability that a particular AR-D ratio is optimal for the given data repetition. More concretely, we can estimate the density of this distribution by sampling from the posterior of the GPR model. The result of this is visualized in the bottom part of Figure 4.
@@ 5.3 Results and discussion
In the first case, it is clearly beneficial to put more weight on the autoregressive training than on masked-diffusion. Yet, training only autoregressively does not lead to any improvement in any experiments within the regular-data region. Even when evaluated purely autoregressively, the differences between $256:0$ and $15:1$ ratios are negligible. Switching to bidirectional evaluation, the single-objective $256:0$ ratio performs poorly while all models trained with ratios between $255:1$ and $15:1$ perform similarly – notably, they all substantially outperform models trained only with masked-diffusion. This is a key finding: even without any data constraint, the dual-objective models achieve stronger masked-diffusion performance than pure masked-diffusion training, despite dedicating only a small fraction of training to the masked-diffusion objective. We hypothesize that the prevalence of the autoregressive objective leads to fast convergence, and that the small amount of masked-diffusion balances its slower convergence by inducing useful modeling priors. This leads us to formulating the first practical recommendation:
Remark 1 (Language modeling under regular data settings).
@@
When training a language model in a regular data setting (16 repetitions or less), train with a small amount of masked-diffusion objective (roughly every 64th sequence) to achieve stronger bidirectional performance than pure masked-diffusion training without losing any autoregressive performance.
@@
In the second data-constrained case, the relation between data repetition, AR-D ratio, and final performance seems more complicated. We risk overfitting by putting too much weight on autoregression and underfitting by focusing too much on masked-diffusion; as evident from Figure 4, the interval of optimal ratios is fairly narrow. On the other hand, the optimal ratios are surprisingly similar for the unidirectional and bidirectional performance. We can notice that the region of optimal ratios is right beneath the region of ratios that lead to overfitting, but the question is how to identify such an AR-D ratio. It is possible to have an alternative interpretation of the ratios and count the number of data repetitions that each objective is individually trained on – then we can see that more than 32 autoregressive repetitions lead to overfitting while fewer than 8 autoregressive repetitions lead to underfitting. Thus, based on the empirical results, our recommendation for this scenario is:
Remark 2 (Data-constrained language modeling).
@@
When training a language model in a data-constrained setting (more than 32 repetitions), choose an autoregressive-diffusion ratio that exposes the autoregressive objective to roughly 16 repetitions of the training data.
@@
Table 2: The normalized autoregressive performance of selected models. We show the results on all nine evaluated tasks for three repetition values; each repetition group contains the results of the best-performing autoregressive-diffusion ratio and of the autoregressive-only model. The scores for each task are normalized so that 0% corresponds to random baseline and 100% is the perfect score. The best result for each dataset size is boldfaced.
@@
   Dual ($63:1$) 05.7 28.6 63.7 35.1 31.1 04.9 17.6 40.9 14.3 26.9
   Autoregressive ($1:0$) 05.9 30.3 61.3 33.5 31.7 03.8 13.6 39.4 15.2 26.1
@@
   Dual ($3:1$) 03.3 28.0 57.9 31.1 26.4 03.6 14.4 36.1 14.6 23.9
   Autoregressive ($1:0$) 05.0 24.9 53.3 28.5 25.4 03.8 09.9 33.3 14.2 22.0
@@
   Dual ($1:7$) 01.7 23.6 56.1 24.8 14.2 01.6 08.5 28.1 13.3 19.1
   Autoregressive ($1:0$)  -1.0 12.3 33.2 06.8 08.1 01.1  -0.5 15.8 08.9 09.4
@@
An obvious question is whether the recommendations hold even at much bigger scale for larger language models. Reliably answering this question would require expensive experimentation, but we believe that the conclusions hold for two reasons. Firstly, according to our results, the optimal AR-D ratios are clearly correlated with overfitting of autoregressive language models. Since the overfitting behavior does not depend on model size according to previous work (NEURIPS2023_9d89448b; prabhudesai2025diffusionbeatsautoregressivedataconstrained), we believe that the optimal AR-D ratios should also not change. Secondly, the relative burden of representing two modes of operation within the learned parameters decreases with model size, so we believe that the benefit of the dual training objective should actually increase with model size.
@@
To put the abstract average scores into another perspective, we look at the individual (normalized) scores per task in Table 2. The results show that the improvement in performance from using a dual objective is observed on a majority of tasks. This is especially true the more repetitions there are. The detailed scores also highlight how effectively the dual objective learns from limited data, reaching nontrivial performance even when exposed to just 256M tokens of training data (under 128 repetitions). We observe similar trends for masked-diffusion evaluation except that as the number of repetitions decreases, the performance gap increases rather than decreases. Detailed performance for the masked-diffusion evaluation can be found in Appendix J.
@@ 5.4 Generalization to prefix language modeling
Prefix language modeling (NEURIPS2019_c20bb2d9; 10.5555/3455716.3455856; pmlr-v162-wang22u) is a promising alternative to the two training objectives investigated in this work. It processes the conditioning part (prefix, $\bm{c}$ in notation from Section 4.1) of a text fully bidirectionally while the completion part ($\bm{w}$ in Section 4.1) is processed autoregressively. Given that our models are trained with both unidirectional and bidirectional attention, we test whether the exposure to both can induce generalization to prefix language modeling without any further training. We repeat the earlier autoregressive evaluation with prefix attention masks and plot the results in Figure 5.
Figure 5: Interpolated prefix results. The figures show the relation between data repetitions (x-axis), autoregressive-diffusion ratios (y-axis), and downstream performance (color-coded). The individual results are interpolated by a GPR model. The right figure demonstrates the relative improvement of prefix-masked evaluation compared to fully unidirectional evaluation (blue color denotes decreased performance and red color denotes a performance increase).
@@
Remark 3 (Induced prefix language modeling).
The autoregressive performance of dual language models can be reliably improved at inference time by processing the conditional part of a prompt fully bidirectionally.
@@ 6 Related work
This paper builds upon the GPT-BERT training objective by charpentier-samuel-2024-bert, validating its effectiveness in a more practical setting. However, there is a long history of papers that tried to combine bidirectional masked language modeling with unidirectional autoregressive modeling: T5 (10.5555/3455716.3455856) and BART (lewis-etal-2020-bart) were the first to train with autoregressive fill-in-the-blank training objectives by relying on encoder-decoder transformer architectures. Later, du-etal-2022-glm proposed GLM, which uses the same objective as T5 while using a simpler decoder-only architecture with a complicated scheme of positional encodings. CM3 by aghajanyan2022cm3 further simplifies training by not requiring any non-standard architectural modifications like the previous work. As they also add autoregressive language-modeling objective, their work is close to our approach – a model trained with CM3 can be used as any other autoregressive model at inference time, similarly to us. However, our objective also generalizes masked-diffusion language modeling and allows for fine-grained balance of the two objectives throughout training. More recently AntLM by yu-etal-2024-antlm proposed to switch from one objective to the other in a curriculum fashion, starting with a short autoregressive training, followed by a long masked language training and finishing on another short autoregressive training. While this does show promise, the transition from one objective to the other leads to forgetting of the previous objective whereas our objective continuously learns both objectives. Other notable works include prefix language models (NEURIPS2019_c20bb2d9; 10.5555/3455716.3455856; pmlr-v162-wang22u) and UL2 (tay2023ul).
@@
Concurrent works by prabhudesai2025diffusionbeatsautoregressivedataconstrained and ni2025difflm have demonstrated that masked-diffusion models outperform autoregressive models in data-constrained training regimes. Our results confirm their findings but we show that using either of these training objectives is never optimal – combining them together is always better, not only in data-constrained settings.
@@
A recent paper by katz-etal-2025-segment shows that using a bidirectional mask on user and system prompts improves performance on a wide variety of tasks, in line with Remark 3. However, for models to be able to use such masks, the authors first need to train adapters. Our work shows that by training both autoregressive and masked-diffusion at the same time, we are able to induce the prefix mask without any additional training.
@@
NEURIPS2023_9d89448b studies the scaling laws of autoregressive models in data-constrained settings with a similar motivation to this paper. They show that autoregressive models cannot meaningfully learn from more than 16 data repetitions – we demonstrate that this value is at least an order of magnitude larger when training with the dual objective.
@@
Our work shares motivation with the autoregressive-diffusion models proposed by 10.5555/3666122.3667859. The diffusion process in that work is biased towards left-to-right denoising, which improved the decoding efficiency of the diffusion language models at that time. Similarly, arriola2025block speeds up decoding of masked-diffusion models by autoregressively generating chunks of tokens where each chunk is decoded by a diffusion process. In both cases, the resulting models are still diffusion models – albeit faster; these approaches do not generalize over autoregressive and masked-diffusion language modeling as our method.
@@
The recent work by xue2025anyorder modifies masked-diffusion language models by parameterizing them with causally-masked transformers, which makes the diffusion models more comparable to standard autoregressive models – decoupling their architectural differences from differences in training objectives. Their conclusion is that masked diffusion alone is a suboptimal objective for language, which is also confirmed by our experiments (Figure 4). However, we found that by simply combining both objectives, we can get the benefits of diffusion without losing any performance.
@@
Large language models are known to reliably follow the empirical scaling laws that describe how their performance should improve with increased compute, model size, and training data. kaplan2020scalinglawsneurallanguage first demonstrated these relationships, showing how the training loss decreases as a power law with respect to these three parameters. These laws were later refined by 10.5555/3600270.3602446, who showed that compute-optimal training requires scaling data and model size together. Related to our work, the scaling laws reveal a fundamental problem: achieving each incremental gain in performance requires exponentially more training data. Thus, data-constrained language modeling is quickly becoming a relevant field of study even for high-resource languages such as English.

### v2

@@ 5.1 Pretraining setup
We train each 470-million-parameter language model (with 360M non-embedding weights) on 32 billion tokens in total. A repetition factor $R$ means we sample a unique subset of size $\nicefrac{{\text{32B}}}{{R}}$ tokens and repeat it $R$ times during training. This total token budget is more than $4\times$ past the Chinchilla compute-optimal point (Hoffmann et al., 2022); we specifically decided to conduct the experiments in this regime as it reflects how modern language models are trained in practice. This compute budget is also large enough to induce non-trivial zero-shot downstream performance, enabling us to measure clear differences between different configurations.
@@
The language models have 24 layers with hidden size of 1 024, their self-attention operations are divided into 16 parallel heads, the feed-forward modules have intermediate size of 3 554, and the vocabulary is set to 51 200 tokens. As for the architecture itself, we follow the usual modifications of the original transformer recipe (Vaswani et al., 2017) – pre-normalization (Nguyen and Salazar, 2019) with RMSNorm (Zhang and Sennrich, 2019), rotational positional embedding (Su et al., 2024) and Swish-gated linear units (Ramachandran et al., 2018; Shazeer, 2020).
@@
The parameters are optimized by the Muon optimizer for faster convergence (Jordan et al., 2024), specifically its variation proposed by Liu et al. (2025). The learning rate is set to 0.007 and decayed according to the warmup-stable-decay (WSD; Hägele et al., 2024) schedule (without warmup steps and 2 048 steps of linear decay). In total, each model is trained for 8 192 steps with 4M tokens in each global batch and with a sequence length of 2 048 tokens. The optimization is regularized by weight decay (with strength of $10^{-1}$) and by an auxiliary z-loss term (with strength of $10^{-4}$; Chowdhery et al., 2022).
@@
Even though we limit the training data to 32B tokens, we deliberately choose a text corpus that is not excessively filtered and that is representative of large-scale web crawls used in practice. We randomly sample English documents with 32B tokens in total from the HPLT v2 corpus (Burchell et al., 2025), which combines extracted webpages from the Internet Archive and CommonCrawl. We also use a smaller disjoint subset to monitor the validation loss. To prevent a potential bias from using an external tokenizer, we train a standard byte-level BPE tokenizer (Gage, 1994) with 51 200 subwords directly on the full training data.
Figure 4: Interpolated unidirectional and bidirectional results. The (a) and (b) figures on top show the relation between repetitions (x-axis) and the autoregressive-diffusion weight $\color[rgb]{0.8,0.0,0.05}\definecolor[named]{pgfstrokecolor}{rgb}{0.8,0.0,0.05}\alpha$} (y-axis); the contours follow the Gaussian process model that interpolates the average performance of language models trained according to the specified settings. The respective results are plotted either as crosses when the model overfitted during training, or as circles. The (c) and (d) figures below visualize the estimated probability that a particular $\color[rgb]{0.8,0.0,0.05}\definecolor[named]{pgfstrokecolor}{rgb}{0.8,0.0,0.05}\alpha$} (y-axis) is optimal for a given number of repetitions (x-axis).
5.2 Searching for the optimal $\color[rgb]{0.8,0.0,0.05}\definecolor[named]{pgfstrokecolor}{rgb}{0.8,0.0,0.05}\alpha$}
We trained and evaluated 50 language models in total, the results are plotted in Figure 4. In order to deal with the noisy nature of this data and to better understand the relation between the amount of data repetitions and the optimal $\color[rgb]{0.8,0.0,0.05}\definecolor[named]{pgfstrokecolor}{rgb}{0.8,0.0,0.05}\alpha$}, we use simple statistical models.
@@
Specifically, we use Gaussian process regression (GPR; Williams and Rasmussen, 1995) with a composite kernel structure to model the relationship between data repetitions, $\color[rgb]{0.8,0.0,0.05}\definecolor[named]{pgfstrokecolor}{rgb}{0.8,0.0,0.05}\alpha$} and downstream performance. The composite kernel consists of a constant kernel multiplied by an anisotropic Matérn kernel ($\nu=1.5$; Stein, 1999) combined additively with a white noise kernel to account for observation noise. The input features are standardized to zero mean and unit variance, and the output features are normalized. The kernel parameters are optimized by L-BFGS-B (Liu and Nocedal, 1989) using SciPy (Virtanen et al., 2020). The resulting interpolations in Figure 4 show regular structure while closely fitting the data with $R^{2}$ over 0.99 in all cases.
@@
The fitted Gaussian process is a probabilistic model of the downstream performance with regard to the amount of data repetition and $\color[rgb]{0.8,0.0,0.05}\definecolor[named]{pgfstrokecolor}{rgb}{0.8,0.0,0.05}\alpha$}. Thus, we can transform this to the probability that a particular $\color[rgb]{0.8,0.0,0.05}\definecolor[named]{pgfstrokecolor}{rgb}{0.8,0.0,0.05}\alpha$} is optimal for the given data repetition. More concretely, we can estimate the density of this distribution by sampling from the posterior of the GPR model. The result of this is visualized in the bottom part of Figure 4.
@@ 5.3 Results and discussion
In the first case, it is clearly beneficial to put more weight on the autoregressive training than on masked-diffusion. Yet, training only autoregressively does not lead to any improvement in any experiments within the regular-data region. Even when evaluated purely autoregressively, the differences between $\color[rgb]{0.8,0.0,0.05}\definecolor[named]{pgfstrokecolor}{rgb}{0.8,0.0,0.05}\alpha$} set to $1$ and $\nicefrac{{15}}{{16}}$ are negligible. Switching to bidirectional evaluation, the single-objective ${\color[rgb]{0.8,0.0,0.05}\definecolor[named]{pgfstrokecolor}{rgb}{0.8,0.0,0.05}\alpha}=1$ performs poorly while all models trained with $\color[rgb]{0.8,0.0,0.05}\definecolor[named]{pgfstrokecolor}{rgb}{0.8,0.0,0.05}\alpha$} values between $\nicefrac{{255}}{{256}}$ and $\nicefrac{{15}}{{16}}$ perform similarly – notably, they all substantially outperform models trained only with masked-diffusion. This is a key finding: even without any data constraint, the dual-objective models achieve stronger masked-diffusion performance than pure masked-diffusion training, despite dedicating only a small fraction of training to the masked-diffusion objective. We hypothesize that the prevalence of the autoregressive objective leads to fast convergence, and that the small amount of masked-diffusion balances its slower convergence by inducing useful modeling priors. This leads us to formulating the first practical recommendation:
@@
Remark 1  Language modeling under regular data settings When training a language model in a regular data setting (16 repetitions or less), train with a small amount of masked-diffusion objective (${\color[rgb]{0.8,0.0,0.05}\definecolor[named]{pgfstrokecolor}{rgb}{0.8,0.0,0.05}\alpha}\approx\nicefrac{{63}}{{64}}$) to achieve stronger bidirectional performance than pure masked-diffusion training without losing any autoregressive performance.
@@
In the second data-constrained case, the relation between data repetition, $\color[rgb]{0.8,0.0,0.05}\definecolor[named]{pgfstrokecolor}{rgb}{0.8,0.0,0.05}\alpha$}, and final performance seems more complicated. We risk overfitting by putting too much weight on autoregression and underfitting by focusing too much on masked-diffusion; as evident from Figure 4, the interval of optimal $\color[rgb]{0.8,0.0,0.05}\definecolor[named]{pgfstrokecolor}{rgb}{0.8,0.0,0.05}\alpha$} values is fairly narrow. On the other hand, the optimal values are surprisingly similar for the unidirectional and bidirectional performance. We can notice that the region of optimal $\color[rgb]{0.8,0.0,0.05}\definecolor[named]{pgfstrokecolor}{rgb}{0.8,0.0,0.05}\alpha$} values is right beneath the region of $\color[rgb]{0.8,0.0,0.05}\definecolor[named]{pgfstrokecolor}{rgb}{0.8,0.0,0.05}\alpha$} values that lead to overfitting, but the question is how to identify such an $\color[rgb]{0.8,0.0,0.05}\definecolor[named]{pgfstrokecolor}{rgb}{0.8,0.0,0.05}\alpha$}. It is possible to have an alternative interpretation of the autoregressive-diffusion weights and count the number of data repetitions that each objective is individually trained on – then we can see that more than 32 autoregressive repetitions lead to overfitting while fewer than 8 autoregressive repetitions lead to underfitting. Thus, based on the empirical results, our recommendation for this scenario is:
@@
Remark 2  Data-constrained language modeling When training a language model in a data-constrained setting (more than 32 repetitions), choose $\color[rgb]{0.8,0.0,0.05}\definecolor[named]{pgfstrokecolor}{rgb}{0.8,0.0,0.05}\alpha$} that exposes the autoregressive objective to roughly 16 repetitions of the training data.
@@
Table 2: The normalized autoregressive performance of selected models. We show the results on all nine evaluated tasks for three repetition values; each repetition group contains the results of the best-performing $\color[rgb]{0.8,0.0,0.05}\definecolor[named]{pgfstrokecolor}{rgb}{0.8,0.0,0.05}\alpha$} and of the autoregressive-only model. The scores for each task are normalized so that 0% corresponds to random baseline and 100% is the perfect score. The best result for each dataset size is boldfaced.
@@
   Dual (${\color[rgb]{0.8,0.0,0.05}\definecolor[named]{pgfstrokecolor}{rgb}{0.8,0.0,0.05}\alpha}=\nicefrac{{63}}{{64}}$) 05.7 28.6 63.7 35.1 31.1 04.9 17.6 40.9 14.3 26.9
   Autoregressive (${\color[rgb]{0.8,0.0,0.05}\definecolor[named]{pgfstrokecolor}{rgb}{0.8,0.0,0.05}\alpha}=1$) 05.9 30.3 61.3 33.5 31.7 03.8 13.6 39.4 15.2 26.1
@@
   Dual (${\color[rgb]{0.8,0.0,0.05}\definecolor[named]{pgfstrokecolor}{rgb}{0.8,0.0,0.05}\alpha}=\nicefrac{{3}}{{4}}$) 03.3 28.0 57.9 31.1 26.4 03.6 14.4 36.1 14.6 23.9
   Autoregressive (${\color[rgb]{0.8,0.0,0.05}\definecolor[named]{pgfstrokecolor}{rgb}{0.8,0.0,0.05}\alpha}=1$) 05.0 24.9 53.3 28.5 25.4 03.8 09.9 33.3 14.2 22.0
@@
   Dual (${\color[rgb]{0.8,0.0,0.05}\definecolor[named]{pgfstrokecolor}{rgb}{0.8,0.0,0.05}\alpha}=\nicefrac{{1}}{{8}}$) 01.7 23.6 56.1 24.8 14.2 01.6 08.5 28.1 13.3 19.1
   Autoregressive (${\color[rgb]{0.8,0.0,0.05}\definecolor[named]{pgfstrokecolor}{rgb}{0.8,0.0,0.05}\alpha}=1$)  -1.0 12.3 33.2 06.8 08.1 01.1  -0.5 15.8 08.9 09.4
@@
An obvious question is whether the recommendations hold even at much bigger scale for larger language models. Reliably answering this question would require expensive experimentation, but we believe that the conclusions hold for two reasons. Firstly, according to our results, the optimal $\color[rgb]{0.8,0.0,0.05}\definecolor[named]{pgfstrokecolor}{rgb}{0.8,0.0,0.05}\alpha$} values are clearly correlated with overfitting of autoregressive language models. Since the overfitting behavior does not depend on model size according to previous work (Muennighoff et al., 2023; Prabhudesai et al., 2025), we believe that the optimal $\color[rgb]{0.8,0.0,0.05}\definecolor[named]{pgfstrokecolor}{rgb}{0.8,0.0,0.05}\alpha$} values should also not change. Secondly, the relative burden of representing two modes of operation within the learned parameters decreases with model size, so we believe that the benefit of the dual training objective should actually increase with model size.
@@
To put the abstract average scores into another perspective, we look at the individual (normalized) scores per task in Table 2. The results show that the improvement in performance from using a dual objective is observed on a majority of tasks. This is especially true the more repetitions there are. The detailed scores also highlight how effectively the dual objective learns from limited data, reaching nontrivial performance even when exposed to just 256M tokens of training data (under 128 repetitions). We observe similar trends for masked-diffusion evaluation except that as the number of repetitions decreases, the performance gap increases rather than decreases. Detailed performance for the masked-diffusion evaluation can be found in Appendix L.
@@ 5.4 Generalization to prefix language modeling
Prefix language modeling (Dong et al., 2019; Raffel et al., 2020; Wang et al., 2022) is a promising alternative to the two training objectives investigated in this work. It processes the conditioning part (prefix, $\bm{c}$ in notation from Section 4.1) of a text fully bidirectionally while the completion part ($\bm{w}$ in Section 4.1) is processed autoregressively. Given that our models are trained with both unidirectional and bidirectional attention, we test whether the exposure to both can induce generalization to prefix language modeling without any further training. We repeat the earlier autoregressive evaluation with prefix attention masks and plot the results in Figure 5.
Figure 5: Interpolated prefix results. The figures show the relation between data repetitions (x-axis), $\color[rgb]{0.8,0.0,0.05}\definecolor[named]{pgfstrokecolor}{rgb}{0.8,0.0,0.05}\alpha$} (y-axis), and downstream performance (color-coded). The individual results are interpolated by a GPR model. The right figure demonstrates the relative improvement of prefix-masked evaluation compared to fully unidirectional evaluation (blue color denotes decreased performance and red color denotes a performance increase).
@@
Remark 3  Induced prefix language modeling The autoregressive performance of dual-objective language models can be reliably improved at inference time by processing the conditional part of a prompt fully bidirectionally.
@@ 6 Related work
This paper builds upon the GPT-BERT training objective by Charpentier and Samuel (2024), validating its effectiveness in a more practical setting. However, there is a long history of papers that tried to combine bidirectional masked language modeling with unidirectional autoregressive modeling: T5 (Raffel et al., 2020) and BART (Lewis et al., 2020) were the first to train with autoregressive fill-in-the-blank training objectives by relying on encoder-decoder transformer architectures. Later, Du et al. (2022) proposed GLM, which uses the same objective as T5 while using a simpler decoder-only architecture with a complicated scheme of positional encodings. CM3 by Aghajanyan et al. (2022) further simplifies training by not requiring any non-standard architectural modifications like the previous work. As they also add autoregressive language-modeling objective, their work is close to our approach – a model trained with CM3 can be used as any other autoregressive model at inference time, similarly to us. However, our objective also generalizes masked-diffusion language modeling and allows for fine-grained balance of the two objectives throughout training. More recently AntLM by Yu et al. (2024) proposed to switch from one objective to the other in a curriculum fashion, starting with a short autoregressive training, followed by a long masked language training and finishing on another short autoregressive training. While this does show promise, the transition from one objective to the other leads to forgetting of the previous objective whereas our objective continuously learns both objectives. Other notable works include prefix language models (Dong et al., 2019; Raffel et al., 2020; Wang et al., 2022) and UL2 (Tay et al., 2023).
@@
Concurrent works by Prabhudesai et al. (2025) and Ni et al. (2025) have demonstrated that masked-diffusion models outperform autoregressive models in data-constrained training regimes. Our results confirm their findings but we show that using either of these training objectives is never optimal – combining them together should always be better, not only in data-constrained settings.
@@
A recent paper by Katz et al. (2025) shows that using a bidirectional mask on user and system prompts improves performance on a wide variety of tasks, in line with Section 5.4. However, for models to be able to use such masks, the authors first need to train adapters. Our work shows that by training both autoregressive and masked-diffusion at the same time, we are able to induce the prefix mask without any additional training.
@@
Muennighoff et al. (2023) studies the scaling laws of autoregressive models in data-constrained settings with a similar motivation to this paper. They show that autoregressive models cannot meaningfully learn from more than 16 data repetitions – we demonstrate that this value is at least an order of magnitude larger when training with the dual objective.
@@
Our work shares motivation with the autoregressive-diffusion models proposed by Wu et al. (2023). The diffusion process in that work is biased towards left-to-right denoising, which improved the decoding efficiency of the diffusion language models at that time. Similarly, Arriola et al. (2025) speeds up decoding of masked-diffusion models by autoregressively generating chunks of tokens where each chunk is decoded by a diffusion process. In both cases, the resulting models are still diffusion models – albeit faster; these approaches do not generalize over autoregressive and masked-diffusion language modeling as our method.
@@
The recent work by Xue et al. (2025) modifies masked-diffusion language models by parameterizing them with causally-masked transformers, which makes the diffusion models more comparable to standard autoregressive models – decoupling their architectural differences from differences in training objectives. Their conclusion is that masked diffusion alone is a suboptimal objective for language, which is also confirmed by our experiments (Figure 4). However, we found that by simply combining both objectives, we can get the benefits of diffusion without losing any performance.
@@
Large language models are known to reliably follow the empirical scaling laws that describe how their performance should improve with increased compute, model size, and training data. Kaplan et al. (2020) first demonstrated these relationships, showing how the training loss decreases as a power law with respect to these three parameters. These laws were later refined by Hoffmann et al. (2022), who showed that compute-optimal training requires scaling data and model size together. Related to our work, the scaling laws reveal a fundamental problem: achieving each incremental gain in performance requires exponentially more training data. Thus, data-constrained language modeling is quickly becoming a relevant field of study even for high-resource languages such as English.